Proceedings of LALSSWF ’98

Paul E. Black pblack@acm.org

1 Introduction

This is a report and proceedings of the 4th Annual LAL SuperScalar Waffle Feed (LALSSWF ’98) [3] held in Provo, Utah, on February 27, 1998. The annual event began three years ago as a bold experiment to apply techniques of modern microprocessor architecture to daily life.^{1} Through arduous research and sometimes acrimonious debate, we decided to experiment with applying the principles of superscalar and pipelined processing to making breakfast. Traditionally held on the last Friday of February, the feed has come to be a landmark in the application of advanced scheduling and processing concepts to feeding starving graduate students, a defensible reason to put off dissertation work, and an asylum for atrocious puns.

2 Attendence

Category  Male Female Total
Professor  2  0  2
from Government  1  0  1
from Industry  1  1  2
Grad Student  3  3  6
earned Ph.D.  3  0  3
earned M.S.  7  4  11
baby  0  1  1
guide dog  1  0  1

Table 1: Demographics of Attendees

Although attendence was off some from last year, the feed still attracted some world-class gastronomes. Table 1 gives a statistical break-down of attendees.

1The real reason was a surprise birthday party for a graduate student who would never have shown up if she knew we knew it was her birthday.

Neither the baby nor the guide dog contributed much, but they were at least as well-behaved as everyone else. The professors contributed adequately, but they were well-behaved, too.

Everybody went away satisfied with the content of the event. Many declared it was one of the most fulfilling assemblages they had ever attended.

3 Papers

The review process for submissions was very demanding this year. In fact, no papers were found to be of adequate quality for presentation.^{2} There weren’t even any puns worth mentioning.

4 Experimental Setup

Due to limited resources^{3}, we used commercially developed material. However we provide the following to allow the reader to repeat the experiment.

• 1 3/4 cup flour

• 2 teaspoons baking powder

• 1/2 teaspoon salt

• 1 tablespoon sugar

• 3 eggs

• 7 tablespoons vegetable oil

• 1 1/2 cups milk

Sift dry ingredients into a bowl. Separate eggs, putting the egg whites in another bowl. Beat egg whites until stiff. Add yokes, oil, and milk to dry ingredients. Beat until smooth. Fold egg whites into batter. Use 1/2 cup of batter per waffle.

Figure 1: Production Curve

Multiple waffle irons allowed superscalar production yielding greater throughput. Speedup is nearly linear, as shown in Figure 1. Batter production, thermochemical refinement, and consumption were pipelined to reduce latency.

2This may be due, in some small part, to the fact that no papers were submitted.

3We were too lazy to start from scratch.

Just as a microprocessor requires many ancillary blocks for bus interface, interrupt processing, virtual memory, etc. [1], the waffles were supported by eggs, sausage, and orange juice.

5 Experimental Results

We applied practical insights^{4} gained from previous years experience to get “improved execution load distribution techniques, improved batter prefetch techniques and some very tasty breakfasts.” [2]

At one point production suffered a stall. We changed the controller and quickly resumed.

6 Panel Discussion

The presence of a guide dog in training sparked a spontaneous panel discussion of other possible applications.

1. Proposal dog, to prepare those tedious and vacuous proposals.

2. Collect signatures dog, to track down people who are never in.

3. Reboot server dog, to go down to the machine room and restart the computer when it hangs.

7 Conclusions

Invite your secretary whenever performing such gustatory conclaves; it helps keep you on their good side. Schedule to begin early; nobody else is sufficiently differently-abled, intellect-wise, to be up that early. Bring a large plastic bag to dispose of any inconvenient evidence.

References

[1] John L. Hennessy and David A. Patterson. Computer Architecture: A Quantitative Approach Morgan Kaufmann, 1990.

[2] LAL Superscalar Waffle Feed ’96. URL: http://lal.cs.byu.edu/people/ jones/lalsswf/96feed.html (accessed: October 20, 1998).

[3] Third Annual LALSSWF Call for Participation. URL: http://lal.cs.byu.edu/ people/bunker/lal/98cfp.html (accessed: October 20, 1998).

4Don’t plug too many waffle irons into the same circuit. Stagger the initial heat up to limit the amperes drawn. Otherwise the building super may wonder why computers are blowing circuit breakers.